OpenStudy (anonymous):

need help plz :) ☆.。.:*・°☆.。.:*・°☆.。.:*・°☆.。.:*・°☆ will fan and medal (⌒.−) A spherical raindrop evaporates at a rate proportial to its surface area . write a differential equation for the volume of the rain drop as a function of time.

OpenStudy (lastdaywork):

First find Volume as a function of time & then make a differential equation by eliminating the arbitrary constant(s)..

OpenStudy (anonymous):

how could i do that ?

OpenStudy (lastdaywork):

"A spherical raindrop evaporates at a rate proportial to its surface area" Can you write the corresponding equation for this ^^ statement ??

OpenStudy (anonymous):

no

OpenStudy (lastdaywork):

Can you tell me the formula for surface area for a sphere ? Please don't let me down :P

OpenStudy (anonymous):

\(\Huge S=4\pi r^2\)

OpenStudy (lastdaywork):

O.O So biggggg.. LOL

OpenStudy (lastdaywork):

BTW, rate of evaporation can be considered as dV/dt (or -dV/dt ; I prefer working with positive, the constant of proportionality can be considered negative). Hence, the sentence "A spherical raindrop evaporates at a rate proportial to its surface area" is equivalent to \[\frac{ dV }{ dt }=kr^2\] where k is a constant.

OpenStudy (lastdaywork):

We already have r as a function of V. k is the only arbitrary constant here..eliminate k to get the final answer..

OpenStudy (anonymous):

plz continue

OpenStudy (lastdaywork):

EOL :P

OpenStudy (anonymous):

plzzzz

OpenStudy (anonymous):

do it or ill tag mashy

OpenStudy (anonymous):

V = volume, t = time, and S = surface area. Then dV dt = −kS, for k > 0.

OpenStudy (anonymous):

it will help u

OpenStudy (anonymous):

\(dv=-ks dt\)?! ot \(dv=-kr^2 dt\)

OpenStudy (anonymous):

ok thx both of you ! @LastDayWork @HARSH123 !